OpenStudy (anonymous):

The wave length(W) of a wave varies inversely as its frequency(F), a wave with a frequency of 1100kHz has a length of 132m. What is the length of a wave with frequency of 600kHz?

OpenStudy (anonymous):

l is prop to 1/f hence vf= constant so l1f1=l2f2 1100 x 132 = l2 x 600 l2 = 244 m

OpenStudy (anonymous):

ty but i dont understand your work will you explain so i can make notes and do the next one

OpenStudy (anonymous):

see if i say that a is inversely prop to b then it means that a = k/b where k is any constant...understand that much?

OpenStudy (anonymous):

yes and inversely means you flip the 2nd equation?

OpenStudy (anonymous):

yes... and as u see a=k/b we can say that ab=k or the product of any a and b is constant..right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so it effectively means that the product of any pair of a and b is equal to the product of any other pair of a and b...get it? like a1 X B1 = K and a2 x b2 = k so we see that a1 x b1 = a2 x b2 right?

OpenStudy (anonymous):

so now do u get it??

OpenStudy (anonymous):

i entered 244m and its wrong the correct answer is 242m

OpenStudy (anonymous):

oh sorry..typo..242

OpenStudy (anonymous):

w= k --- f 132= k ---- 1100 k=145,200 w=145,200 ------ f w=145,200 ------ 600 w= 242m

OpenStudy (anonymous):

yeah...right

OpenStudy (anonymous):

ty